OpenStudy (anonymous):

x=(1/x)y^2 The directrix of the parabola is: x=-2 x=2 y=-2

OpenStudy (jdoe0001):

\(\large x = \frac{1}{x}y^2 \ \ ?\)

OpenStudy (anonymous):

yeahh

OpenStudy (jdoe0001):

\(x = \cfrac{1}{x}y^2 \implies x^2 = y^2 \implies x^2-y^2=0\) maybe it's me, but you have an \(x^2\) and a \(y^2 \) both "x" and "y" terms are squared, which means is not a parabola it'd be an ellipse if both were positive, one negative would make it a hyperbola, and they'd equal to 1, which in this case, they do not so can't say, but doesn't look like any conic equation

OpenStudy (anonymous):

thats exactlyy what i thoughtt... thankss

OpenStudy (jdoe0001):

yw